OpenStudy (anonymous):

In his pocket Ted has $1 bill and some quarters. Let q represent the number of quaters

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero

hero (hero):

@w.timora, which choice do you think is correct?

OpenStudy (anonymous):

@hero the last one

hero (hero):

How much is a quarter worth?

OpenStudy (anonymous):

25 cent

hero (hero):

How do you write 25 cent in decimal form?

OpenStudy (anonymous):

0.25?

hero (hero):

Exactly

OpenStudy (anonymous):

ok so so it has to be one answer with 0.25

OpenStudy (anonymous):

how do I figure that out?

hero (hero):

Read the question once more, then think about it.

OpenStudy (anonymous):

is it the second one?

hero (hero):

Yes. The next time you come across a question like this, what you should do is use actual values to test out the equation. For example, suppose Ted had 1 dollar and 1 quarter. If that's the case then he should have $1.25 in his pocket. You think 1 + 0.25q is correct so test it: Let q = 1, then 1 + 0.25q = 1 + 0.25(1) = 1 + 0.25 = 1.25 Suppose he had 1 dollar and 2 quarters in his pocket. Then he should have $1.50. Testing that: 1 + 0.25q = 1 + 0.25(2) = 1 + 0.50 = 1.50 If you keep going, you'd realize option B is the only option that gives the correct result every time.

OpenStudy (anonymous):

@hero Thank you so much!